Bradley owns a school supplies store. For inventory, he found out that he sold a total of 250 notebooks worth P5,050. A small notebook costs P15 while a big one is P25. How many big notebooks were sold?

Thank you !!! :D :D :D

Answer by JulietG(1812)

Total = 5050
If all of the notebooks sold were big notebooks, the total would be 250*25, or 6250.
6250-5050=1200
The difference in the cost between the big notebooks and the small notebooks is 25-15, or 10.
1200/10 = 120
120 of the notebooks were small. That leaves 130 big notebooks sold.
120 * 15 = 1800
130 * 25 = 3250
1800+3250 = 5050